OpenStudy (anonymous):

what are the two values of x that satisfy log_x 2=log_3 X

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

Use the definition of the log like we did with the others here.

OpenStudy (anonymous):

\[2 = log_3x \iff\ ?\]

OpenStudy (anonymous):

x=3^(log_x2) idk....there are two xs and it's seems ambiguous....like, if i do 2=x^(log_3 X) then, it's just weird..

OpenStudy (anonymous):

I'm not sure there are 2 values for x that apply here though.

OpenStudy (anonymous):

Wait, where's the 2 x's

OpenStudy (anonymous):

I only see one.

OpenStudy (anonymous):

log_x 2=log_3 X

OpenStudy (anonymous):

oh, that's different.

OpenStudy (anonymous):

But even so, the rule still applies.

OpenStudy (anonymous):

\[Let\ k = log_3 x \implies 3^k = x\] But that means that \[k = log_x 2 \implies x^k = 2\]

OpenStudy (anonymous):

So if we take the first equation and raise it to the power of k we have \[3^{kk} = x^k = 2\]

OpenStudy (anonymous):

i get your previous answer. But why are you raising them to the pwr k??

OpenStudy (anonymous):

So that I can set them equal.

OpenStudy (anonymous):

oh~

OpenStudy (anonymous):

Because now we can take the ln of both side and have \[k^2(ln 3) = (ln 2)\]

OpenStudy (anonymous):

so k = ?

OpenStudy (anonymous):

the sqroot of ln2/ln3. But how did you go from x^k=2 to ln2?

OpenStudy (anonymous):

I didn't. I went from \[3^{k^2} = 2 \implies k^2(ln\ 3) = (ln\ 2)\]

OpenStudy (anonymous):

oh, ok. got it.

OpenStudy (anonymous):

Also \[k = \pm \sqrt{(\frac{ln 2}{ln 3})}\] Which is where you get your two solutions from.

OpenStudy (anonymous):

Though you still have to solve for x.

OpenStudy (anonymous):

wait, so there's two values for x too?

OpenStudy (anonymous):

Yes, because there are 2 values for k, and k is related to x.

OpenStudy (anonymous):

\[k = log_3 x\] Remeber?

OpenStudy (anonymous):

haha right~

OpenStudy (anonymous):

So \[x = 3^k\]

OpenStudy (anonymous):

So the two values that satisfy that equation for x are?

OpenStudy (anonymous):

3^sqroot ln3/ln2 and the other one is the same except it's the negative root of ln3/ln2. yes? (please say yes ^^:)

OpenStudy (anonymous):

It's ln2/ln3, but otherwise yes.

OpenStudy (anonymous):

YAY <3